An amount of money is invested at 8% interest per year, compounded quarterly. After a number of years, the


accumulated amount in the account is R7 365,00. The investment earned R2 000,00 interest in this period.


If the accumulated amount is left in the account (the interest rate remains the same) for another period that


is one year longer than the first period, the accumulated amount in the account will then be


[1] R10 312,78


[2] R29 363,56


[3] R10 944,00


[4] R10 699,8

1
Expert's answer
2022-04-18T17:04:20-0400

7365 - 2000 = 5365


"5365 \cdot(1+0.02)^{4x} = 7365"

"(1.02)^{4x} = 7365\/5365"

"4x = 16"

x = 4

4 years


Because of it one year longer than the first period, so 4+1 = 5 years

"5365 \cdot(1+0.02)^{4\cdot 5} = 10944"

Answer: C